LynnMathena:

What is the solution? {4x+3y=543x+9y=108

xxemilianaxx:

Can you send a ss the problem doesn't really make any sense

mhanifa:

You can solve it by elmination

xxemilianaxx:

How would you know, You can't even tell what the problem is

mhanifa:

3 times the first equation minus the second one

mhanifa:

@xxemilianaxx wrote:
How would you know, You can't even tell what the problem is
This is a system of equations

xxemilianaxx:

If your problem is 4x+3y=54| 3x+9y=108 that's what I'm guessing then.. 4 x + 3 y = 54 / · ( - 3 ) 3 x + 9 y = 108 - 12 x - 9 y = - 162 3 x + 9 y = 108 - 9 x = - 54 | x = - 54 : ( - 9 ) = 6 4 · 6 + 3 y = 54 24 + 3 y = 54 3 y = 54 - 24 3 y = 30 | y = 30 : 3 = 10 Thus, ( 6, 10 ) Since we know thats what ( x , y ) is ( 6 , 10 )